How many months would it take to earn $450 on a deposit of $8,000 earning 4.5% simple interest?
Answer by addingup(3677)

Calculate time: t = I/Pr where t is the time, I is the interest amount, P the principal, and r the rate of interest:
t = 450/(8000*0.045) = 1.25 years